# A New Algorithm for Zero-Sum Linear-Quadratic Stochastic Differential Games in Infinite Horizons

### v1

@@ 3.3 Convergence Analysis of Iterative Sequence
4 Numerical Experiments
To verify the effectiveness of the proposed algorithm, we present a specific numerical example and randomly generate the system parameters as follows:
@@ 4 Numerical Experiments
After 13 outer iterations, the algorithm converges to the stabilizing solution of the problem within the specified error tolerance. The norm of the equation residual is $4.9409\times 10^{-6}$. The specific values of the stabilizing solution to GTARE (3) (retained to six decimal places) are presented as follows:
$\begin{bmatrix}4.490400&0.330564&-1.002100\\ 0.330564&4.895330&-0.781543\\ -1.002100&-0.781543&8.616318\end{bmatrix}$.
@@
Figure 3: This simulation diagram clearly illustrates the dynamic process of a multivariable system under the balanced control strategies of two players.
Figure 1, 2 depicts the evolution of the real parts of eigenvalues across different matrices with external iterations (k=1to 13)(k=1\,to\,13). And the both Figure show a consistent pattern: the real parts of the eigenvalues of $Z^{(k)}$ and $M_{(k)}$ respectively decrease over iterations, approaching 0 by the sixth iteration and continuing to drop, notably featuring a drastic change in the initial $M_{(1)}$ eigenvalues due to one-time disturbances from the opposing player. They collectively reflect an interesting characteristic: When the set $\mathcal{A}$ is allowed to be non-empty, Player 2 starts from a dominant position and injects controls to stabilize the system. Meanwhile, the opposing Player 1 is dissatisfied with the current state. In response, Player 2 makes tentative concessions by continuously granting Player 1 certain advantages. This leads to a gradual reduction in Player 1’s dissatisfaction, which can be intuitively observed from Figure 1,2. Through continuous adjustments, the system approaches its equilibrium position.
Figure 3 shows that the system’s three state variables ($x_{1}$, $x_{2}$, $x_{3}$) undergo a rapid and intense transient process at the initial moment, with the instantaneous pulse of $x_{3}$ being particularly prominent. However, this potential instability trend is quickly suppressed, and all states converge within an extremely short time to reach stability near the zero point. The control input diagrams in the middle and below reveal the reason for this phenomenon: the control commands of the two players exhibit strong, oppositely directed antagonistic pulses in the initial phase. Their interaction does not disrupt the system; instead, it collectively offsets the system’s inherent unstable dynamics, ultimately driving the states toward equilibrium. This is consistent with the conclusions derived from the theoretical research in the Remark 2.6.3 of [28].
Together, these figures verify the effectiveness of our algorithm and indirectly reveal the mechanism by which the algorithm iteratively finds the equilibrium solution.

### v2

@@ 3.3 Convergence Analysis of Iterative Sequence
$)\displaystyle+\sum_{l=1}^{r}(C_{l}+D_{l,1}K_{1}(\tilde{P})+D_{l,2}K_{2}(\tilde{P}))^{\top}(P^{*}-\tilde{P})(C_{l}+D_{l,1}K_{1}(\tilde{P})+D_{l,2}K_{2}(\tilde{P}$))
[…]
From Proposition 3.7, we have $P^{*}\succeq\tilde{P}$. Combining this with $\tilde{P}_{\tilde{L}}=\tilde{P}$, we obtain $\tilde{P}\preceq P^{*}\preceq\tilde{P}_{\tilde{L}}=\tilde{P}$. Therefore, $P^{*}=\tilde{P}$ is the unique stabilizing solution to the stochastic GTARE (3). ∎
4 Numerical Example
To verify the effectiveness of the proposed computational method, we present a specific numerical example and randomly generate the system parameters as follows:
@@ 4 Numerical Example
After 13 external iterations, the iterative sequence converges to the stabilizing solution of the problem within the specified error tolerance. The norm of the equation residual is $4.9409\times 10^{-6}$. The specific values of the stabilizing solution to the stochastic GTARE (3) (retained to five decimal places) are presented as follows:
$\begin{bmatrix}4.49040&0.33056&-1.00210\\ 0.33056&4.89533&-0.78154\\ -1.00210&-0.78154&8.61632\end{bmatrix}$.
@@
Figure 3: This simulation diagram clearly illustrates the dynamic process of the system state under the balanced control strategies of two players.
Fig. 1 and Fig. 2 illustrate the evolution of the real parts of eigenvalues across different matrices over external iterations $(k=1,\dotsb,13)$. Both figures reveal a consistent trend: the real parts of the eigenvalues of $Z^{(k)}$ and $M_{(k)}$ respectively decrease with iterations, approaching 0 by the sixth iteration and continuing to decline thereafter. Notably, the eigenvalues of $M_{(1)}$ undergo a drastic change, attributed to one-time disturbances from the opposing player. These collectively reflect an interesting characteristic: when the set $\mathcal{A}$ is non-empty, Player 2 initiates from a dominant position and applies controls to stabilize the system. Meanwhile, Player 1 is dissatisfied with the current values of the performance functional. In response, Player 2 makes tentative concessions by continuously granting Player 1 certain advantages, leading to a gradual reduction in Player 1’s dissatisfaction. Through such continuous adjustments, the performance functional gradually approaches its equilibrium values.
Fig. 3 illustrates the dynamic evolution of the system’s three state variables ($x_{1}$, $x_{2}$, $x_{3}$) starting from a randomly generated initial state. Initially, the system displays pronounced unstable dynamics, evidenced by the rapid and intense transients of all states—with $x_{3}$ exhibiting a particularly large initial deviation. This unstable trend is quickly suppressed, however, and the states converge asymptotically to the origin. The control input diagrams in the middle and lower panels explain this phenomenon: the controls exerted by the two players involve strong, oppositely directed antagonistic injections. Within the stability injection set $\mathcal{K}$, the two players pursue mutually acceptable equilibrium control strategies to ensure the system’s equilibrium across all directions. While their control injection are antagonistic, their interaction yields a combined effect that counteracts the system’s inherent unstable dynamics, collectively driving the states toward equilibrium. Notably, despite the zero-sum game framework, the two parties demonstrate a cooperative tendency in maintaining system stability. These observations align with the theoretical findings presented in Remark 2.6.3 of [Sun2020_book].
Together, these figures verify the effectiveness of our computational method and indirectly reveal the mechanism by which the method iteratively finds the equilibrium solution.